HUBBLE SPACE TELESCOPE DAILY REPORT #2162 PERIOD COVERED: 0000Z (UTC) 07/09/98 - 0000Z (UTC) 07/10/98 Daily Status Report as of 191/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in yesterday's report and in HSTAR 6575, the reacquisition for the second iteration of this proposal failed and the take data flag went down. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/1 7861 (A Search for Extrasolar Giant Planets in the Nearby TW Hya Association) The Near Infrared Camera (Camera One) was used to search for extrasolar giant planets on the nearby star USNO-21 (a.k.a., TW Hya). One of the primary goals of NASA's mission is the identification of formed and nascent planetary systems. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed NIC/1/2 7856 (Jets, Cones and the Alignment Effect in High-Z Radio Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the radio galaxy 3C324. Two distinct physical mechanisms -- anisotropic quasar illumination and jet-cloud interactions -- are required to explain the general properties of high-z {z > 0.5} radio galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxy IR14378-36. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.8 Completed STIS/CCD/MA1/MA2 7931 (Scattered Light in the Echelle Modes) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to study scattered light in the STIS Echelle Modes by making observations of the O class star CPD59D260. The effect is strongest at the shortest red wavelengths. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.9 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.10 Completed NIC/1 7821 (SINS: The Supernova Intensive Study - Cycle 7+: Scarlet SINS) The Near Infrared Camera (Camera One) was used to make observations of the super nova LMC-SN1987 which is located in the Large Magellanic Cloud. Supernovae create the chemical history of the Universe, energize the interstellar gas, and stand at the center of efforts to understand the extragalactic distance scale. The observations were completed as scheduled, and no problems were noted. 1.11 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.12 Completed WF/PC-2 6776 (A Comparative HST Imaging Study of Radio Galaxies and Hosts) A clear understanding of the nature of the host galaxies of radio-quiet quasars, radio-loud quasars and radio galaxies is of central importance in any attempt to unify and relate the three main classes of powerful active galaxies. Existing HST studies of quasar hosts, whilst producing many interesting new results, have yielded surprisingly little hard information on the properties of the hosts themselves. This imaging study of 3C59 using the WF/PC-2 will remedy this situation by undertaking redder, line-free observations of our three matched AGN samples. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: The SSR EDAC error counter was cleared three times using ROP SR-1. As described in HSTAR 6576, the STIS MCE-2 reset at 190/1303Z while the voltage was on. The MAMA-2 was recovered nominally via SMS commanding at 190/1651Z when the high voltage was turned on. The DF-224 SPC memory was dumped to CCS strings A and B at 191/0542Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH